OpenStudy (anonymous):

what is the soultion to the equation 6x+8=-4

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

6x + 8 = -4 subtract 8 from both sides 6x = -4 -8 divide both sides by 6 x = -12/6 x = -2